OpenStudy (anonymous):

you have a 35 min. communte to work. Since you leave early, the trip going to work is easier than the trip home. You can travel to work in the same time that it takes you to make it 28 miles back home. Your average speed coming home is 5 MPH slower than your average speed going to work. What is your average speed going to work?

OpenStudy (anonymous):

I got 15.5...?

OpenStudy (anonymous):

15.5 MPH

OpenStudy (chris):

hmmm. let's see. So on the way to work: 1. x * 35min/(60min/hour) = 28miles

OpenStudy (anonymous):

why does that work?

OpenStudy (chris):

hmm, or are they not the same length?

OpenStudy (anonymous):

They are the same length, I'm sure.

OpenStudy (chris):

Hmmm, but in that case it doesn't make as much sense.

OpenStudy (anonymous):

It seems like it would be?

OpenStudy (chris):

If the trips are the same length (28 miles), and you know how long it takes to get to work (35 min), you would know the average speed

OpenStudy (chris):

by distance/time

OpenStudy (chris):

I think what the problem is trying to do is give you three unknowns, and enough information for three equations

OpenStudy (chris):

three unknowns would be x=speed to work, y = speed back from work, d= distance to work

OpenStudy (anonymous):

Gotcha.

OpenStudy (anonymous):

I did 1/35 + 1/28 = 1/t(unknown time)

OpenStudy (anonymous):

I found the LCD to be 980t.

OpenStudy (chris):

ah ok, here we go: 1. so y*35/60 = 28 (with y being the speed coming back from work) 2. x = y+5 (speed going to work is five mph faster) so now just solve for y in the first equation and add 5

OpenStudy (chris):

sorry for the confusion heh. So that means that y=48mph, and x = 53mph